OpenStudy (anonymous):

What is the vertex of the graph of y = –4(x + 2)2 + 5? (2, 5) (–2, 5) (5, –2) (5, 2)

OpenStudy (anonymous):

This is vertex form. y=a(x-h)^2+k (h,k) is the vertex. Can you see what it would be?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

You can always just look at it and tell the vertex when its in that form. compare y=a(x-h)^2+k to y=-4(x+2)^2+5 h would be -2 and k would be 5 right? so vertex is (-2,5)